In five years a woman will be three times as old as her daughter will be. Three years ago, the woman was five times as old as her daughter was then. Find their present ages.
Answer by stanbon(75887)

In five years a woman will be three times as old as her daughter will be. Three years ago, the woman was five times as old as her daughter was then. Find their present ages.
-------------------
Equations:
w+5 = 3(d+5)
w-3 = 5(d-3)
------------------
Rearrange::
w - 3d = 10
w - 5d = -12
---------------------
Subtract and solve for "d":
2d = 22
Daughter's age is 11 yrs now
---------------------------
Solve for "w":
w - 3d = 10
w - 33 = 10
Woman's age is 43 yrs now
---------------------
Cheers,
Stan H.
-------------